OpenStudy (anonymous):

A car weighing 1600 kr travels at 90 km/h. The driver brakes and after 10s the car is standing still. How far did the car travel after the driver began braking?

OpenStudy (anonymous):

So far I've got an average deacceleration of 2,5m/s, but how do I get from there to the distance traveled?

OpenStudy (anonymous):

u can use the suvat equations:

OpenStudy (anonymous):

justwait

OpenStudy (anonymous):

suvat equations: v=u+at missing s s=1/2(u+v)t missing a s=ut+1/2at^2 missing v s=vt-1/2at^2 missing u v^2=u^2+2as missing t

OpenStudy (anonymous):

s=distance u=initial velocity (speed) v=final speed a=acceleration or deceleration t=time

OpenStudy (anonymous):

i don't understand what this part means "A car weighing 1600 kr " is it the weight of the car?

OpenStudy (anonymous):

yes, sorry that's a typo, it's supposed to be kg

OpenStudy (anonymous):

thanks for your reply, I'll just stare at this for a while :)

OpenStudy (anonymous):

so u have got s=? u=90 km/h in physics u have to convert units to m/s v=0 m/s a=? t=10 s

OpenStudy (anonymous):

can u calculate the u in m/s?

OpenStudy (anonymous):

@math0101

OpenStudy (anonymous):

u can calculate that then choose one of the suvat equations

OpenStudy (anonymous):

u will probably use this equatiion: s=1/2(u+v)t

OpenStudy (anonymous):

then substitute the values and then u have the distance travelled by the car

OpenStudy (anonymous):

I've realized I need to look into this further so I'm looking through the chapter on constant acceleration now that I know what I need to know / understand. Thanks for your help :)

OpenStudy (anonymous):

answer is 125 metres

OpenStudy (anonymous):

it's ok pleasure @math0101

OpenStudy (anonymous):

u r right @rexparkar but u don't need to give answers like this but let others try first especially the ones who asked the question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its a constant retardation problem